An object with a mass of 2.3 x 102 g has a leftward force of 6.2 N applied to it.  What is the resulting acceleration of the object?


1
Expert's answer
2021-05-11T16:35:02-0400

We can find the resulting acceleration of the object from the Newton's Second Law of Motion:


"F=ma,""a=\dfrac{F}{m}=\dfrac{6.2\ N\ [leftward]}{0.23\ kg}=27\ \dfrac{m}{s^2}\ [leftward]."